Add the following fractions:

1/5 + 5/4 6/2 + 3/4
1
Expert's answer
2011-02-14T04:54:22-0500
1/5+5/4 = (4*1 + 5*5)/20 = 29/20
6/2+3/4 = (2*6+3)/4 = 15/4